OpenStudy (anonymous):

Please help! This problem has me nine different kinds of confused! Water freezes at 0 degrees Celsius and 32 degrees Fahrenheit. Water boils at 100 degrees Celsius and 212 degrees Fahrenheit. a. Write degrees Celsius as a linear function of degrees Fahrenheit. b. what is the slope of your linear equation? What does it mean? c. what is the y-intercept and what does it mean?

OpenStudy (solomonzelman):

I only know that C=273+K, I took physics long time ago, don't know....

OpenStudy (anonymous):

Let's talk about A. For every one degree in Fahrenheit, how many degrees are there in Celsius?

OpenStudy (anonymous):

well, when water is at 0 degrees celsius, fahrenheit is at 32 degrees. So, for every 32 degrees fahrenheit, is there a degree celsius? Or am I really off with that guess?

OpenStudy (anonymous):

Hmm. I am sorry. I made an error. Let's start from scratch and derive the equation from scratch :)

OpenStudy (anonymous):

The first thing we are going to do is make a graph.

OpenStudy (anonymous):



OpenStudy (anonymous):

Here I have graphed the points where water both freezes and boils at these temperatures.

OpenStudy (anonymous):

Are you following?

OpenStudy (phi):

you are given two points (32,0) and (212,100) where the first number (the "x" is Fahrenheit) and y is Celsius) now find the equation of a line, given that you know 2 points. one way: find the slope then use the point-slope formula y - y0 = m(x - x0)